# Hot Rocks Survey III: A deep eclipse for LHS 1140c and a new Gaussian process method to account for correlated noise in individual pixels

### v1

@@ 4.1.1 Primary reduction: Aperture extraction analysis
We performed a joint-fit of the three eclipses using priors on the central (primary) transit time , system scale , impact parameter , and radius ratio from the NIRISS/SOSS transit observation of LHS 1140c as analysed in Cadieux et al. (2024b). As the NIRISS/SOSS analysis assumed a circular orbit with fixed period, we used priors for the period and eccentricity parameters and from the joint-fit of radial velocity and previous transit observations (performed before the JWST transit observation) in Cadieux et al. (2024a). Including the JWST transit observation from Cadieux et al. (2024b) within a joint-fit of the RV data could help to tighten these constraints further but is beyond the scope of this paper. Table 11 shows a summary of these priors. As the time of primary transit is tightly constrained with this prior, the central eclipse time is largely determined by the eccentricity parameters (see Appendix C).
@@
where is the timestamp of the first integration being fit and the midpoint of the observation, is the baseline flux out-of-(secondary) transit, is a linear slope in baseline flux, is the amplitude of the settling ramp and is the lengthscale of the ramp. When fitting for a combined exponential ramp and linear slope, there were issues with degeneracies between long exponential ramps with large values and the linear slope , so in practice we also subtract off the tangent line to the exponential ramp at the last integration which helped to reduce this degeneracy when was large.
@@
where is the height scale and the length scale of correlated noise and is the amplitude of white noise.
@@ 4.1.2 Primary reduction: Pixel-fitting analysis
For the pixel-fitting method, aperture extraction was not needed and the reduced data was used directly. The same eclipse model and priors were used as for the aperture extraction method.
Figure 9 shows the choice of pixels within 5px of the PSF centre for the pixel-fitting method overlaid over the median frame of the PSF. It is also shown how it compares to the 5px circular aperture used for aperture extraction. Note that unlike aperture extraction, pixel-fitting can weight away from lower signal-to-noise pixels similar to optimal extraction, so we can choose to include more pixels to try improve our constraints. However, including more pixels becomes more computationally expensive particularly as there are many more parameters to recover, and the change in constraints was found to be quite negligible when pixels 8.5px away from the centre were included (covered in Appendix G).
@@ 4.2 Secondary data reduction
We perform a secondary reduction using a completely independent pipeline, which has been previously used in August et al. (2025) and will be described in more detail in the upcoming paper Gressier et al., in prep. This pipeline employs the transitspectroscopy (Espinoza 2022) wrapper for the initial stages of the jwst pipeline, skipping the reference pixel correction and incorporating a custom jump correction. The identified 10 Hz heater noise is corrected. Once the groups are converted to slopes, we proceed to Stage 2. We compute a median frame, which is then used for the following step. To estimate the background shape, we mask pixels within a box centered around the source. The masked pixels are interpolated using a linear interpolation with neighboring pixels. This frame is then smoothed using a Gaussian filter with a standard deviation of 7 pixels. The resulting median background image is subtracted from all frames. Pixel light curves are extracted, and outliers are identified using a median filter. These outliers are replaced with the median value of each pixel. Once we obtain a background corrected, cleaned image, we proceed to flux extraction.
@@ 4.2.1 Secondary reduction analysis
We use juliet (Espinoza et al. 2019) to fit the extracted light curves from this secondary reduction. The three observations are jointly fitted using five detrending methods: a linear slope in time (L), a single exponential (E), a linear slope with a single exponential (L+E), a linear slope with a Gaussian process (L+GP), and a linear slope with a single exponential and a Gaussian process (L+E+GP). We use the Matérn 3/2 kernel, which is described by:
@@ 4.3.1 Results from primary reduction analysis
The results from joint-fitting the light curves are shown in Table 4. While the results are all highly consistent, the pixel-fitting method tends to be a bit more conservative with larger uncertainties - which is similar to the simulations where there were only flux-conserving systematics present (i.e. the FCS Scenario, see Section 3.5). Note that pixel-fitting while only accounting for flux-conserving systematics performs almost identically to aperture extraction without a GP, so any differences in the size of uncertainties are likely arising purely from the constraints on systematics which are not flux-conserved.
@@
45 | 32782 | 21575 | 27279 | 26845 | |
60 | 32185 | 22680 | 24388 | 26249 |
We also examined fitting for the eclipse depth individually for each dataset. In this case, we assumed a circular orbit and used constraints on the central eclipse time from the other two eclipses as a prior on each individual fit as each eclipse alone gave poor constraints on the eclipse time. These priors are listed in Table 10. We experimented with varying the initial time being cut () and the effect of fitting for an exponential ramp or using a GP. Due to the computational cost of the pixel-fitting method, we only tested these variations with aperture extraction. The results are presented in Table 5 showing the eclipse depths individually for each eclipse as well as the combined result from the three eclipses calculated using error propagation (i.e. not a joint-fit).
The individual eclipse depths show that fitting for an exponential ramp has a large effect on the result. Given the strong settling ramp seen at the beginning of the first eclipse (see Figure 8), the eclipse depth may be biased if this ramp is not accounted for - even after cutting a full hour at the start. This would explain why the linear model without a GP (denoted ’L’ in the Model column) recovers very deep eclipses (¿400ppm) inconsistent with the other two eclipses. The linear model with a GP (L+GP) performs slightly differently and recovers large uncertainties likely because the GP height scale is increasing to fit this initial settling ramp. As initial settling effects are not in line with a stationary Gaussian process, we would prefer not to use a GP to fit them and so our preferred choice of model is the linear with exponential ramp and a GP (L+E+GP) model. This performs very similarly to the linear with exponential ramp without a GP (L+E) model but including a GP is slightly more conservative as it can account for additional systematics which could potentially lie in these data. We see that both these models give eclipse depths that are all within of each other and varying the time cut between 30-60 minutes has minimal impact on the result.
@@ 4.3.2 Results from secondary reduction analysis
The secondary analysis, conducted using an independent pipeline and the juliet package to fit the light curve, yields consistent results. Table 6 presents the eclipse depth results for various detrending methods in the combined fit. The combined eclipse depths remain consistent within 1 of those from the primary analysis, which used classic aperture photometry or pixel-fitting, except for the L and L+E models in the secondary analysis. However, these models are strongly disfavored, compared to the other models. The use of Gaussian Processes inflates the uncertainty.
@@ 4.5 Variations in detector settling
We investigated potential explanations for the variations in detector settling seen at the start of our three LHS 1140c observations (shown in Figure 8). It was noticed that the previous MIRI observation for the first eclipse was a calibration observation which was taking darks for MIRI/Imaging. This observation also shows a strong negative slope very different from the start of the other two eclipses. Note that this calibration observation ended ¿5 hours before the start of the first LHS 1140 observation (the previous observation of any kind was using NIRSpec/IFU). However, MIRI will typically leave the previous filter it used in position until almost immediately before the next MIRI observation begins. In this case, the OPAQUE filter was kept in place until switching to the F1500W filter, leaving a long time for each pixel to settle without receiving flux. For the other two eclipses, the F1000W and F770W filters were left in place until right before the observation.
@@
Note that there is no shutter on MIRI - the detector is continuously observing the sky and constantly resetting after each read - this means that the telescope will spend time observing LHS 1140 with the previous filter in the time after slewing to the target but before switching filter. It has already been noted that this change in filters can lead to persistence effects from different amounts of background flux in each filter (Dicken et al. 2024). However, the PSF of LHS 1140 observed through the different wavelength filters may also have sufficient time to produce persistence effects and affect the beginning of our observations. Examining this using the JWST engineering mnemonics for telescope orientation and MIRI filter position appears to support this possibility, with the telescope pointing in a similar direction to the observation period for about 10-15 minutes before the filter is switched to F1500W (see Figure 11).
@@
To test this idea, each of the observations in the Hot Rocks survey which used the SUB256 subarray were reduced in the same manner as described in Section 4.1. The four observations in the survey which used either the SUB64 or SUB128 subarrays would require a different background subtraction and occur on completely separate pixels on the detector so we restricted our comparison to just the SUB256 observations. After we had performed the same background subtraction and 5px radius aperture extraction as for the LHS 1140c eclipses, we took the first 30 minutes of each observation (clipping the first integration to avoid the ”first exposure” effect) and fit a best-fit line to it. We then compare the slope of the first 30 minutes of each observation to the previous filter used. The previous filter used was found with the IMIR_HK_FW_CUR_POS engineering mnemonic, which provides the MIRI filter currently in position. In all cases, the observations are being switched to the F1500W filter. See Figure 12 for a plot of these results.
@@
There appears to be a pattern where the observations which previously used short wavelength filters (i.e. F560W, F770W, F1000W) show flat or positive slopes at the start, while longer wavelength filters up to the F2000W filter show increasingly negative slopes. The F2550W filter doesn’t quite match this trend but still shows a negative slope. The OPAQUE filter performs similarly to the long wavelength filters with a strong negative slope, while observations which previously had the P750L prism used for MIRI/LRS have slightly negative or flat slopes. The fourth eclipse of L 231-32b was the only observation which previously used a coronagraphic filter (F1550C), with this observation being consistent with a flat slope. Note that observations which both used the same previous filter tend to have quite consistent slopes, even when the targets being observed are different. This suggests that it may be possible to improve the consistency of the initial detector settling by switching filter earlier (e.g. before slewing or before guide star acquisition), which might enable better characterisation of detector settling and allow us to clip less data.
@@
It is worth noting that if we avoid performing any background subtraction we recover similar slopes in the aperture extracted light curves. Most settling slopes are shifted by less than 0.01%/hr if a background subtraction is not performed, with the most affected observation being the first LHS 1140c eclipse which is shifted 0.035%/hr flatter (i.e. hr to /hr). This implies that any trend is primarily driven by the strong settling effects seen in the central pixels of the PSF and not related to persistence effects in the background pixels. The strong settling of the central pixels is visualised in Figure 13, which shows the pixel light curves centred on the PSF for the initial 30 minutes of each LHS 1140 observation. Consider that when the F770W filter is in position prior to the third observation, this filter should produce a times narrower PSF which has about seven times greater total flux compared to the F1500W filter (based on the filter bandpasses and BT-Settl models). This combination of a narrower PSF and a bandpass which collects more flux could likely explain the significant differences in the persistence effects seen on the central pixels compared to an observation which previously had the OPAQUE filter in and hence was exposed to very little light in the time immediately before the observation began. One caveat is that there is some time required to switch filter (s per filter position; Wright et al. 2015), so the filters which will be in-between the previous filter and the F1500W filter may also introduce some persistence effects, although there is significantly less time for this to occur.
@@ 4.6 Pixel-fitting parameters recovered
It is notable that the systematics parameters are all consistent between each dataset. None of the datasets detect common systematics (with height scale ) or find strong evidence of independent pixel systematics from the shared heightscale (). The amplitude of the flux-conserving systematics () - which may be interpreted as the amplitude of the PSF changing shape over time - appears to be well constrained. The correlations in noise arising from interpixel capacitance in the horizontal (), vertical () and diagonal () directions are also tightly constrained and greater than zero. Combining these constraints through error propagation we get %, % and % - consistent with expected values (Morrison et al. 2023).
The length scale in time of the systematics is also well constrained. However, there was a prior placed on each parameter taken from the posterior of the joint-fit accounting for common systematics but not independent pixel systematics. This was because the addition of independent pixel systematics could create a degeneracy between very gradual systematics in each pixel and the slope in each pixel being fit for in the mean function. As a result, the constraints on are more closely related to the timescale of the flux-conserving systematics and so could represent the rate at which the PSF is changing shape (in units of days).
Finally, the parameters describing correlations in background scatter do not definitively constrain correlations in background noise either across all pixels () or along the rows () or columns (). The inclusion of fitting for correlations in noise along each row and column was largely a conservative choice as there are a number of row/column effects seen with MIRI and the lack of a detection does not rule these systematics out but instead puts an upper bound on their presence in these data. However, the background subtraction is not perfect so almost certainly should be non-zero and all three eclipses do appear to favour the value being at least somewhat larger than the minimum prior bound of DN/s. By including more pixels or if our datasets had more integrations, this might provide the method more information to constrain the remaining level of background scatter in the data. This could be interesting as a way to compare different background subtraction strategies and identify which method best minimises background scatter.
@@ 4.7 Absolute flux calibration
As our eclipse depth measurements correspond to the planet-to-star flux ratio , our measurement of the dayside planetary flux is sensitive to our models of flux from the host star across the F1500W filter bandpass. In addition, stellar models are required to determine the amount of radiation the dayside is exposed to, which will determine the expected flux emitted for each model. Inaccuracies in our stellar models may therefore affect our eclipse spectra models in two different ways. While we do not have spectra to examine the accuracy of these models across near to mid-infrared, we can use our observations to perform an absolute flux calibration and examine how closely the BT-Settl models match the observed flux across the F1500W bandpass.
@@
We performed the absolute flux calibration following the procedure used for the JWST Absolute Flux Calibration program (Gordon et al. 2025). This involved running the standard JWST calibration pipeline but skipping the EMI correction step emicorr and turning off the after-jump flagging and shower detection steps of the jump step. We skipped the photom step in stage 2 and instead used the values given in Gordon et al. (2025) to convert from DN/s to millijanskys (mJy). photutils was used for the centroiding and aperture extraction, with an aperture radius of px and background annulus radii of 8.63px and 11.45px - matching the radii used for the program. A conversion factor is used to convert the aperture extracted flux using a finite aperture to the total flux expected from an ”infinite” aperture and this factor also accounts for the oversubtraction of background flux due to some of the PSF being included within the background annulus.
@@
We performed the flux calibration for each integration of our observations and calculated the mean flux. This resulted in flux values of mJy, mJy and mJy for each eclipse respectively, with the uncertainties determined using the reported uncertainty in the calibration factor of (the uncertainties due to Poisson noise are negligible in comparison). In comparison, if we calculate the expected flux weighted across the F1500W bandpass using a grid of BT-Settl (CIFIST) models (Allard et al. 2011), we find an expected flux of mJy. The uncertainty in this value propagates uncertainties in the stellar distance pc from Gaia DR3 (Gaia Collaboration et al. 2021) and in the stellar radius , stellar effective temperature and surface gravity from Cadieux et al. (2024a).
@@
This makes the measured flux larger than the flux expected from the BT-Settl models after marginalising over and - or greater than the interpolated BT Settl model with and fixed to mean values444One caveat is that as we are scaling the BT-Settl models based on the stellar radius and system distance, it is possible that the BT-Settl models are accurate but for example the stellar radius value from Cadieux et al. (2024a) is inaccurate.. If this increased stellar flux is real and only deviates from the BT-Settl models in a region around the F1500W bandpass, then we may expect this to roughly shift our bare rock and atmospheric eclipse models to be shallower by affecting the ratio. In principle, this should not dramatically affect our conclusions given our uncertainty in the measured eclipse depth (and in fact may provide stronger evidence on the lack of an atmosphere). However, this does not account for how stellar model inaccuracies may affect the radiative transfer/energy balance equations, or how the BT-Settl models may deviate from the stellar spectrum at other wavelengths. We must therefore advise some caution in the interpretation of our eclipse models. We expect that future observations of M-dwarf spectra at near to mid-infrared will help us improve the accuracy of these models as there are currently significant challenges facing the modeling their spectra (Shields et al. 2016).
@@
The absolute flux value can be used in combination with the eclipse depth to calculate the brightness temperature of LHS 1140c. This is performed by multiplying the eclipse depth, absolute flux value and the square of the ratio of the distance the system is away to the radius of the planet. This is compared to integrating Planck’s law over the F1500W bandpass and finding a brightness temperature which corresponds to this level of emission. We performed this by calculating the emission from Planck’s law over a range of brightness temperatures between 200K and 800K in 0.1K intervals and interpolating to invert emission to brightness temperature. Marginalising over the uncertainty in the eclipse depth from aperture extraction with a GP ( ppm) as well as the system distance and planetary radius, we recover a brightness temperature of K. This is close to the maximum brightness temperature we might expect from a (smooth) zero albedo bare rock with no heat redistibution of K, while full heat redistibution even with zero albedo would be inconsistent with this value at only K (and higher albedos would be even colder). However, this is only a simple calculation which does not take into account effects such as the Greenhouse effect, scattering or thermal beaming which are discussed in Section 5. This calculation does have the benefit of not requiring the use of (potentially unreliable) stellar models while the atmospheric modelling in Section 5 requires the use of stellar models to account for the wavelength-dependent effects of absorption/emission and scattering.

### v2

@@ 4.1.1 Primary reduction: Aperture extraction analysis
We performed a joint fit of the three eclipses using priors on the central (primary) transit time, ; system scale, ; impact parameter, ; and radius ratio, , from the NIRISS/SOSS transit observation of LHS 1140c, as analysed in Cadieux et al. (2024b). As the NIRISS/SOSS analysis assumed a circular orbit with fixed period, we used priors for the period, , and eccentricity parameters, and , from the joint fit of radial velocity and previous transit observations (performed before the JWST transit observation) in Cadieux et al. (2024a). Including the JWST transit observation from Cadieux et al. (2024b) within a joint fit of the RV data could help to tighten these constraints further but is beyond the scope of this paper. Table 10 shows a summary of these priors. As the time of primary transit is tightly constrained with this prior, the central eclipse time is largely determined by the eccentricity parameters (see Appendix C).
@@
where is the time stamp of the first integration being fit and the midpoint of the observation, is the baseline flux out-of-(secondary) transit, is a linear slope in baseline flux, is the amplitude of the settling ramp and is the lengthscale of the ramp. When fitting for a combined exponential ramp and linear slope, there were issues with degeneracies between long exponential ramps with large values and the linear slope ; thus, in practice, we also subtracted the tangent line to the exponential ramp at the last integration, which helped reduce this degeneracy when was large.
@@
where is the height scale, the length scale of correlated noise, and is the amplitude of white noise.
@@ 4.1.2 Primary reduction: Pixel-fitting analysis
For the pixel-fitting method, the aperture extraction was not needed and the reduced data were used directly. The same eclipse model and priors were used as for the aperture extraction method.
Figure 9 shows the choice of pixels within 5px of the PSF centre for the pixel-fitting method overlaid over the median frame of the PSF. It is also shown how it compares to the 5px circular aperture used for aperture extraction. We note that unlike aperture extraction, pixel-fitting can weight away from lower S/N pixels similar to optimal extraction, so we can choose to include more pixels to try improve our constraints. However, including more pixels becomes more computationally expensive particularly as there are many more parameters to recover, and the change in constraints was found to be quite negligible when pixels 8.5px away from the centre were included (covered in Appendix G).
@@ 4.2 Secondary data reduction
We performed a secondary reduction using a completely independent pipeline, which has been previously used in August et al. (2025) and will be described in more detail in the upcoming paper (Gressier et al., in prep.). This pipeline employs the transitspectroscopy (Espinoza 2022) wrapper for the initial stages of the jwst pipeline, skipping the reference pixel correction and incorporating a custom jump correction. The identified 10 Hz heater noise is corrected. Once the groups are converted to slopes, we proceed to Stage 2. We compute a median frame, which is then used for the following step. To estimate the background shape, we mask pixels within a box centered around the source. The masked pixels are interpolated using a linear interpolation with neighboring pixels. This frame is then smoothed using a Gaussian filter with a standard deviation of 7 pixels. The resulting median background image is subtracted from all frames. Pixel light curves are extracted, and outliers are identified using a median filter. These outliers are replaced with the median value of each pixel. Once we obtain a background corrected, cleaned image, we proceed to flux extraction.
@@ 4.2.1 Secondary reduction analysis
We use juliet (Espinoza et al. 2019) to fit the extracted light curves from this secondary reduction. The three observations are jointly fitted using five detrending methods: a linear slope in time (L), a single exponential (E), a linear slope with a single exponential (L+E), a linear slope with a GP (L+GP), and a linear slope with a single exponential and a GP (L+E+GP). We use the Matérn 3/2 kernel, which is described by:
@@ 4.3.1 Results from primary reduction analysis
The results from joint-fitting the light curves are shown in Table 4. While the results are all highly consistent, the pixel-fitting method tends to be a bit more conservative with larger uncertainties - which is similar to the simulations where there were only flux-conserving systematics present (i.e. FCS scenario; see Section 3.5). We note that pixel-fitting while only accounting for flux-conserving systematics performs almost identically to aperture extraction without a GP, so any differences in the size of uncertainties are likely arising purely from the constraints on systematics which are not flux-conserved.
@@
| 45 | 32782 | 21575 | 27279 | 26845 | |
| 60 | 32185 | 22680 | 24388 | 26249 |
We also tested fitting for the eclipse depth individually for each dataset. In this case, we assumed a circular orbit and used constraints on the central eclipse time from the other two eclipses as a prior on each individual fit as each eclipse alone gave poor constraints on the eclipse time. These priors are listed in Table 10. We experimented with varying the initial time being cut () and the effect of fitting for an exponential ramp or using a GP. Due to the computational cost of the pixel-fitting method, we only tested these variations with aperture extraction. The results are presented in Table 4 showing the eclipse depths individually for each eclipse as well as the combined result from the three eclipses calculated using error propagation (i.e. not a joint-fit).
The individual eclipse depths show that fitting for an exponential ramp has a large effect on the result. Given the strong settling ramp seen at the beginning of the first eclipse (see Figure 8), the eclipse depth may be biased if this ramp is not accounted for - even after cutting a full hour at the start. This would explain why the linear model without a GP (denoted ‘L’ in the Model column) recovers very deep eclipses (¿400ppm) inconsistent with the other two eclipses. The linear model with a GP (L+GP) performs slightly differently and recovers large uncertainties likely because the GP height scale is increasing to fit this initial settling ramp. As initial settling effects are not in line with a stationary Gaussian process, we would prefer not to use a GP to fit them and so our preferred choice of model is the linear with exponential ramp and a GP (L+E+GP) model. This performs very similarly to the linear with exponential ramp without a GP (L+E) model but including a GP is slightly more conservative as it can account for additional systematics which could potentially lie in these data. We see that both these models give eclipse depths that are all within of each other and varying the time cut between 30-60 minutes has minimal impact on the result.
@@ 4.3.2 Results from secondary reduction analysis
The secondary analysis, conducted using an independent pipeline and the juliet package to fit the light curve, yields consistent results. Table 5 presents the eclipse depth results for various detrending methods in the combined fit. The combined eclipse depths remain consistent within 1 of those from the primary analysis, which used classic aperture photometry or pixel-fitting, except for the L and L+E models in the secondary analysis. However, these models are strongly disfavored, compared to the other models. The use of GPs inflates the uncertainty.
@@ 4.5 Variations in detector settling
We investigated potential explanations for the variations in detector settling seen at the start of our three LHS 1140c observations (shown in Figure 8). It was noticed that the previous MIRI observation for the first eclipse was a calibration observation which was taking darks for MIRI/Imaging. This observation also shows a strong negative slope very different from the start of the other two eclipses. We note that this calibration observation ended ¿5 hours before the start of the first LHS 1140 observation (the last observation of any kind was carried out using NIRSpec/IFU). However, MIRI typically leaves the previous filter it used in position until almost immediately before the next MIRI observation begins. In this case, the OPAQUE filter was kept in place until switching to the F1500W filter, leaving a long time for each pixel to settle without receiving flux. For the other two eclipses, the F1000W and F770W filters were left in place until right before the observation.
@@
We note that there is no shutter on MIRI, as the detector is continuously observing the sky and constantly resetting after each read; this means that the telescope will spend time observing LHS 1140 with the previous filter in the time after slewing to the target but before switching filter. It has already been noted that this change in filters can lead to persistence effects from different amounts of background flux in each filter (Dicken et al. 2024). However, the PSF of LHS 1140 observed through the different wavelength filters may also have sufficient time to produce persistence effects and affect the beginning of our observations. Examining this using the JWST engineering mnemonics for telescope orientation and MIRI filter position appears to support this possibility, with the telescope pointing in a similar direction to the observation period for about 10-15 minutes before the filter is switched to F1500W (see Figure 11).
@@
To test this idea, each of the observations in the Hot Rocks survey which used the SUB256 subarray were reduced in the same manner as described in Section 4.1. The four observations in the survey which used either the SUB64 or SUB128 subarrays would require a different background subtraction and occur on completely separate pixels on the detector so we restricted our comparison to just the SUB256 observations. After we had performed the same background subtraction and 5px radius aperture extraction as for the LHS 1140c eclipses, we took the first 30 minutes of each observation (clipping the first integration to avoid the ‘first exposure’ effect) and fit a best-fit line to it. We then compare the slope of the first 30 minutes of each observation to the previous filter used. The previous filter used was found with the IMIR_HK_FW_CUR_POS engineering mnemonic, which provides the MIRI filter currently in position. In all cases, the observations are being switched to the F1500W filter. See Figure 12 for a plot of these results.
@@
There appears to be a pattern where the observations which previously used short wavelength filters (i.e. F560W, F770W, F1000W) show flat or positive slopes at the start, while longer wavelength filters up to the F2000W filter show increasingly negative slopes. The F2550W filter doesn’t quite match this trend but still shows a negative slope. The OPAQUE filter performs similarly to the long wavelength filters with a strong negative slope, while observations which previously had the P750L prism used for MIRI/LRS have slightly negative or flat slopes. The fourth eclipse of L 231-32b was the only observation which previously used a coronagraphic filter (F1550C), with this observation being consistent with a flat slope. We note that observations which both used the same previous filter tend to have quite consistent slopes, even when the targets being observed are different. This suggests that it may be possible to improve the consistency of the initial detector settling by switching filter earlier (e.g. before slewing or before guide star acquisition), which might enable better characterisation of detector settling and allow us to clip less data.
@@
It is worth noting that if we avoid performing any background subtraction we recover similar slopes in the aperture extracted light curves. Most settling slopes are shifted by less than 0.01%/hr if a background subtraction is not performed, with the most affected observation being the first LHS 1140c eclipse which is shifted 0.035%/hr flatter (i.e. hr to /hr). This implies that any trend is primarily driven by the strong settling effects seen in the central pixels of the PSF and not related to persistence effects in the background pixels. The strong settling of the central pixels is visualised in Figure 13, which shows the pixel light curves centred on the PSF for the initial 30 minutes of each LHS 1140 observation. Consider that when the F770W filter is in position prior to the third observation, this filter should produce a times narrower PSF which has about seven times greater total flux compared to the F1500W filter (based on the filter bandpasses and BT-Settl models). This combination of a narrower PSF and a bandpass which collects more flux could likely explain the significant differences in the persistence effects seen on the central pixels compared to an observation which previously had the OPAQUE filter in and hence was exposed to very little light in the time immediately before the observation began. One caveat is that there is some time required to switch filter (s per filter position; Wright et al. 2015), so the filters which are in between the previous filter and the F1500W filter could also introduce some persistence effects, although there is significantly less time for this to occur.
@@ 4.6 Pixel-fitting parameters recovered
It is notable that the systematics parameters are all consistent between each dataset. None of the datasets constrain the presence of common systematics (with height scale ) or found strong evidence of independent pixel systematics from the shared heightscale (). The amplitude of the flux-conserving systematics () – which could be interpreted as the amplitude of the PSF changing shape over time – appears to be well constrained. The correlations in noise arising from interpixel capacitance in the horizontal (), vertical (), and diagonal () directions are also tightly constrained and greater than zero. Combining these constraints through error propagation we get %, %, and % - consistent with expected values (Morrison et al. 2023).
The length scale in time of the systematics, , is also well constrained. However, there was a prior placed on each parameter taken from the posterior of the joint-fit accounting for common systematics but not independent pixel systematics. This was because the addition of independent pixel systematics could create a degeneracy between very gradual systematics in each pixel and the slope in each pixel being fit for in the mean function. As a result, the constraints on are more closely related to the timescale of the flux-conserving systematics and so could represent the rate at which the PSF is changing shape (in units of days).
Finally, the parameters describing correlations in background scatter do not definitively constrain correlations in background noise either across all pixels () or along the rows () or columns (). The inclusion of fitting for correlations in noise along each row and column was largely a conservative choice as there are a number of row/column effects seen with MIRI and the lack of a detection does not rule these systematics out but instead puts an upper bound on their presence in these data. However, the background subtraction is not perfect so almost certainly should be non-zero and all three eclipses do appear to favour the value being at least somewhat larger than the minimum prior bound of DN/s. By including more pixels or if our datasets had more integrations, this might provide the method more information to constrain the remaining level of background scatter in the data. This could be interesting as a way to compare different background subtraction strategies and identify which method best minimises the background scatter.
@@ 4.7 Absolute flux calibration
As our eclipse depth measurements correspond to the planet-to-star flux ratio, , our measurement of the dayside planetary flux is sensitive to our models of flux from the host star across the F1500W filter bandpass. In addition, stellar models are required to determine the amount of radiation the dayside is exposed to, which will determine the expected flux emitted for each model. Inaccuracies in our stellar models may therefore affect our eclipse spectra models in two different ways. While we do not have spectra to examine the accuracy of these models across near to mid-infrared, we can use our observations to perform an absolute flux calibration and examine how closely the BT-Settl models match the observed flux across the F1500W bandpass.
@@
We performed the absolute flux calibration following the procedure used for the JWST Absolute Flux Calibration program (Gordon et al. 2025). This involved running the standard JWST calibration pipeline but skipping the EMI correction step emicorr and turning off the after-jump flagging and shower detection steps of the jump step. We skipped the photom step in stage 2 and instead used the values given in Gordon et al. (2025) to convert from DN/s to millijanskys (mJy). photutils was used for the centroiding and aperture extraction, with an aperture radius of px and background annulus radii of 8.63px and 11.45px - matching the radii used for the program. A conversion factor is used to convert the aperture extracted flux using a finite aperture to the total flux expected from an ‘infinite’ aperture and this factor also accounts for the oversubtraction of background flux due to some of the PSF being included within the background annulus.
@@
We performed the flux calibration for each integration of our observations and calculated the mean flux. This resulted in flux values of mJy, mJy, and mJy for each eclipse respectively, with the uncertainties determined using the reported uncertainty in the calibration factor of (the uncertainties due to Poisson noise are negligible in comparison). In comparison, when we calculate the expected flux weighted across the F1500W bandpass using a grid of BT-Settl (CIFIST) models (Allard et al. 2011), we find an expected flux of mJy. The uncertainty in this value propagates uncertainties in the stellar distance pc from Gaia DR3 (Gaia Collaboration et al. 2021) and in the stellar radius, , stellar effective temperature, , and surface gravity, , from Cadieux et al. (2024a).
@@
This makes the measured flux larger than the flux expected from the BT-Settl models after marginalising over and - or greater than the interpolated BT Settl model with and fixed to mean values666One caveat is that as we are scaling the BT-Settl models based on the stellar radius and system distance, it is possible that the BT-Settl models are accurate but for example the stellar radius value from Cadieux et al. (2024a) is inaccurate.. If this increased stellar flux is real and only deviates from the BT-Settl models in a region around the F1500W bandpass, then we may expect this to roughly shift our bare rock and atmospheric eclipse models to be shallower by affecting the ratio. In principle, this should not dramatically affect our conclusions given our uncertainty in the measured eclipse depth (and in fact may provide stronger evidence on the lack of an atmosphere). However, this does not account for how stellar model inaccuracies may affect the radiative transfer and energy balance equations or how the BT-Settl models may deviate from the stellar spectrum at other wavelengths. We must therefore advise some caution in the interpretation of our eclipse models. We expect that future observations of M-dwarf spectra at near to mid-infrared will help us improve the accuracy of these models as there are currently significant challenges facing the modeling their spectra (Shields et al. 2016).
@@
The absolute flux value can be used in combination with the eclipse depth to calculate the brightness temperature of LHS 1140c. This is performed by multiplying the eclipse depth, absolute flux value and the square of the ratio of the distance the system is away to the radius of the planet. This is compared to integrating Planck’s law over the F1500W bandpass and finding a brightness temperature which corresponds to this level of emission. We performed this by calculating the emission from Planck’s law over a range of brightness temperatures between 200K and 800K in 0.1K intervals and interpolating to invert emission to brightness temperature. Marginalising over the uncertainty in the eclipse depth from aperture extraction with a GP ( ppm) as well as the system distance and planetary radius, we recover a brightness temperature of K. This is close to the maximum brightness temperature we might expect from a (smooth) zero albedo bare rock with no heat redistibution of K, while full heat redistibution even with zero albedo would be inconsistent with this value at only K (and higher albedos would be even colder). However, this is only a simple calculation which does not take into account effects such as the greenhouse effect, scattering, or thermal beaming (discussed in Section 5). This calculation does have the benefit of not requiring the use of (potentially unreliable) stellar models while the atmospheric modelling described in Section 5 requires the use of stellar models to account for the wavelength-dependent effects of absorption and emission as well as scattering.